OpenStudy (anonymous):

A quadratic equation is shown below: x2 + 5x + 4 = 0 Part A: Describe the solution(s) to the equation by just determining radicand. Show your work. (3 points) Part B: Solve 4x2 -12x + 5 = 0 using an appropriate method. Show the steps of your work, and explain why you chose the method used. (4 points) Part C: Solve 2x2 -10x + 3 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points)

OpenStudy (anonymous):

B 4x2 - 12x + 5 = 0 4x2 - 2x -10x +5 = 0 2x(2x -1) - 5(2x -1)= 0 (2x-5) (2x-1) =0 2x-5=0 2x-1=0 2x=5 2x=1 x=5/2 x=1/2

OpenStudy (anonymous):

So that is the splitting method right?

OpenStudy (anonymous):

factorization

OpenStudy (anonymous):

2x2-10x+3=0 a=2 b=10 c= 3

OpenStudy (anonymous):

now u can solve it by quadratic formula square root b2 - 4ac

OpenStudy (anonymous):

And how exactly would you do that?

OpenStudy (anonymous):

I understand the first part and group factoring i just dont understand the quadratic formula square root thing he said

OpenStudy (anonymous):

And im still kind of confused about Part A and how to tell what the solutions are just by looking at a radicand?

OpenStudy (campbell_st):

the quadratic formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] it can be used for a quadratic in the form \[ax^2 + bx + c = 0\]

OpenStudy (campbell_st):

part A is asking you to use the discriminant \[\Delta = b^2 - 4ac\] to determine the type of solutions you will have 2 unequal real rational roots, 2 unequal real irrational roots or no real roots... which means the solutions are complex numbers

OpenStudy (anonymous):

What exactly is a radicand?

OpenStudy (campbell_st):

radicand is the part inside a square root... in this case it is referring to the general quadratic formula... it is more commonly known as the discriminant